• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M.Shabestari

    Log entry time 08:17:24 on May 4,2008

    Entry number 236875

    keyword=Owl Shift Summary

    Owl shift summary May 4, 2008

    SL : Mitra Shabestari
    TO : Mitra Shabestari
    3rd : Stefano Marrone


    Summary
    ========

    pi0 Production : Kinematics F Part (III)

    BigBite = 48 degrees, BigBite current = 390.0 A (setpoint); 388.1 A (readback)
    LHRS = 14.5 degrees, LHRS momentum = 0.997 GeV/c
    RHRS = 105.0 degrees, RHRS momentum = 0.4596 GeV/c

    Target: 6 cm LH2 with collimator in loop1
    Beam current: 3.5 uA

    ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

    00:00 Shift started.
    Run# 4690 (LHRS), and run# 20862 (RHRS) have already started.
    After these runs we are going to change the target to Al.

    00:25 ----> First Al run ,Run# 4691.Beam current = 3.5 uA.
    Stopped the run to change the prescales. Deadtime is just a few percent......events = 293 K
    00:31 ----> Run# 20863 ; started Right HRS.....events = 76 K

    00:36 ----> Run# 4692 ; events = 1 M

    00:57 ----> Run# 4693 ; events = 1 M

    01:15 ----> Finished two 1 M runs with Al. Moved target to LH2 with collimator

    01:29 ----> Run# 4694 ; First production run with "LH2 with collimator" target.
    HV on S1 and S14 tripped (@ 01:31),so we stopped the run.....events = 349 K

    01:38 ----> Run# 4695 ; events = 1 M

    01:38 ----> Run# 20864 ; started Right HRS.....events = 144 K

    01:50 ----> Run# 4696 ; events = 1 M

    02:06 ----> Run# 4697 ; events = 1 M

    02:19 ----> Run# 4698 ; events = 1 M

    02:33 ----> Run# 4699 ; events = 1 M

    02:56 ----> Run# 4700 ; events = 577 K

    03:08 ----> MCC called. They have problem with RF, and beam will be down for a short time.

    ---------- Moved target to empty. I called MCC to find out about beam. They are waiting for technicians.

    06:31 ----> MCC called. They are going to send the beam.
    I checked the beam positions; they are fine. Moved target to LH2 with collimator.
    We start taking production data.

    06:46 ----> Run# 20865 ; started Right HRS.....events =

    06:46 ----> Run# 4701 ; events = 1 M

    07:03 ----> Run# 4702 ; events = 1 M

    07:16 ----> Run# 4703 ; events = 1 M

    07:29 ----> Run# 4704 ; events = 1 M

    07:43 ----> Run# 4705 ; events = 1 M

    07:55 ----> Run# 4706 ; events = 1 M

    08:09 ----> Run# 4707



    ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

    No beam from 3:08 till 6:31.
    After getting beam back we started taking production data, which went fine till the end of the owl shift.